# A background for thermal photons in heavy ion collisions

### v1

@@ I Introduction
The major sources of photons in the A-A or p-p collision systems include hard scattering, secondary hard scattering, soft processes, initial state and final state radiations, and resonance decays [12]. The weak showers can also be a source of photons but their effects are very small. Other exotic processes involving one extra dimension can also produce photons via graviton exchange. However, such processes are beyond the scope of this work [13]. The major distinction between a p-p collision and an A-A collision is the existence of QGP which itself emits photons, such photons are called thermal photons. Many theoretical approaches have been made to study direct photons like the elliptic fireball, thermalized glasma, hydrodynamic simulations of the “fireball” evolution, etc [14, 15, 16, 17]. The recent experiments under the RHIC Beam Energy Scan (BES) measured the direct photon spectra for lower collision energies like $\sqrt{s_{NN}}=$ 62.4 GeV and 39.0 GeV. The measurements found the excess of direct photons which are thermal. Thermal photons can also be found in smaller collision systems like d-Au and 3He-Au. These systems show QGP-like signals such as jet quenching and elliptic flow but fail to produce signals like baryon enhancement. The existence of thermal photons in these systems would imply the existence of QGP [18, 19]. However, effectively distinguishing thermal photons from other sources is a difficult problem due to the contribution of a large number of processes. The major aim of this study is to find the contributions from different processes in photon spectra which will help to distinguish thermal photons from others. We are using PYTHIA8 Angantyr for this study [20, 21]. The Angantyr model combines individual NN sub-collisions from PYTHIA8 based on the FRITIOF model. The model does not consider a hot thermalized medium (QGP). Hence, it can give a clear picture of the non-collective photon spectra in various collisions. A detailed description of the model can be found in the next section.
@@ II Angantyr
PYTHIA8 recently developed the Angantyr algorithm to simulate nucleus-nucleus collisions and proton-nucleus collisions [20, 21]. The FRITIOF program and the wounded nucleon model inspired the model with added effects from partonic hard scattering [22]. The positions and the number of the interacting nucleons and binary nucleon-nucleon collisions are calculated using Glauber-model-based eikonal approximation in impact-parameter space and Gribov’s corrections on color fluctuations are implemented to include diffractive excitation [23, 24], which appears due to the fluctuations in the nucleon substructure [20]. The effects of such fluctuations can be found in the corresponding paper [25]. The hard partonic sub-collisions, normalized by nucleon-nucleon sub-collisions, play a major role at high energies. The model considers two interaction situations between the projectile and target nucleons. The first scenario treats the collisions as pp-like non-diffractive (ND) processes. In the second situation, a wounded projectile nucleon may interact with several target nucleons, resulting in secondary non-diffractive and diffractive sub-collisions. The interaction likelihood between wounded nucleons is characterized as elastic, nondiffractive, secondary nondiffractive, single-diffractive, or double-diffractive. All subevents are then combined to simulate a p-A or A-A event. This can give a better and more accurate description of a heavy ion event than a scaled p-p event[26].
@@
Figure 3: The photon spectra in Au Au collisions at 200 GeV with rescattering.
@@
Where $\sqrt{s_{ref}}$ is the reference center of mass(c.m) energy. The multiplicity of the produced particles is sensitive to $p^{ref}_{T0}$ [30, 27]. A lower value of $p^{ref}_{T0}$ can lead to high multiplicity. The value of $p^{ref}_{T0}$ is tuned for various collision systems and energies tuned by matching the charged particle multiplicity $\frac{dN_{ch}}{d\eta}$ with the experimental data. The final hadronization is done using the LUND string fragmentation model and the probability of creating hadrons from the initial state of partons is computed using LUND area law [31]. The channels for prompt photon production used for the work are $gg\rightarrow q\gamma,q\bar{q}\rightarrow g\gamma$ $,gg\rightarrow g\gamma$ $,q\bar{q}\rightarrow\gamma\gamma$ , and $gg\rightarrow\gamma\gamma$, the matrix elements are from the corresponding papers [32, 33, 34]. These processes are also included in the soft part of the model and their production is controlled using $p^{0}_{T}$.
Besides the hard processes, the Quantum Electrodynamics (QED) processes during Parton showers are a major source of photons. The charged particles produce photons in the initial state showers by Bremmstraulung. QED shower in PYTHIA 8 incorporates a fully coherent multipole treatment of photon radiation off systems of charged fermions, vectors, and scalar particles, as well as photon splittings to pairs of charged fermions [35, 36, 37, 38, 39]. The photon fluxes from the heavy ions are modeled according to the corresponding works[40, 41].
@@ III.1 Different sources of photons
The photon spectra can be distinguished in two parts: 1. Direct photons and 2. Decay photons. In Fig.1 we have shown the photon spectra for central Au-Au collisions at 200 GeV, the direct and decay photons are shown separately for better visualization. It can be seen that the yield from the decay is much higher in magnitude than that of the prompt photons. The major contributor to the decay photon yield is the neutral pion ($\pi^{0}$) followed by $\eta$ meson, while the contributions from $\eta^{{}^{\prime}}$ and $\omega$ are small but still greater than the prompt photons. The other particles such as $\Lambda$, $\phi$, etc. which are not removed from the data in the experiments have a little contribution compared to other sources but they exceed the prompt photons at low $p_{T}$ and their contributions should be considered for a more accurate identification of direct photons.
The prompt photons are a combination of primary hard scattering, Multiparton interactions (MPI), and parton showers. The Angantyr simulations with all processes enabled, with MPI=off, and with both MPI=off and ISR=off are shown in Fig 1. This suggests that apart from primary hard scattering photons are produced in the secondary interactions and Bremmstrahulang-like processes. It can be seen that most of the low $p_{T}$ photons are produced from the initial state radiation (ISR) in the early stage of collision. The MPIs increase the low $p_{T}$ yield slightly due to the contribution of the soft processes, while it decreases the high $p_{T}$ photons. This can be attributed to the interactions of photons with other hadrons. In PYTHIA8, the soft part of interactions with photons is modeled according to the Vector Meson Dominance (VMD) model. The structure of the photon in the PYTHIA8 framework includes VMD and Generalised Vector Meson Dominance (GVMD), where the photon briefly changes to a vector meson (or an excited vector meson) while interacting with a hadron. Regardless of the configuration the model matches with the experimental data [28] at high $p_{T}$ ($>5$ GeV) within error bars, this is expected as the high $p_{T}$ photons are due to the contribution from primary hard scatterings. The prompt photon yield in Angantyr is lower than a typical scaled p-p collision yield. This is due to separating sub-collisions into ND, DD, and SD interactions. A clear depiction can be seen in Fig 1.
@@
The processes after the hadronisation like the hadronic rescatterings also produce photons. The rescattering is significantly high in a heavy-ion event compared to a typical p-p event due to its high multiplicity, hence it should be considered while studying the final state particles. The physics processes are similar to low-energy QCD processes [21]. The hard QCD processes and MPIs are not included as the energy of these collisions will be way below 10 GeV. A detailed rescattering framework used in this work can be found in the corresponding paper[43]. The rescattering starts early in Angantyr due to the absence of QGP, hence its effects in Angantyr are more pronounced than the QGP-based models like UrQMD and SMASH which use rescattering as afterburner[44, 45]. Considering the low energy scales of the processes the prompt photon production is unlikely but the same can not be told about the decay photons. Many exotic resonances are formed during the rescatterings and decay into photons (these photons are often considered direct but we will refer to them as decay photons due to the nature of their origin). The major contributors are particles like the $f_{0}(600)$ meson which is produced abundantly in rescattering due to its low mass. In a rescattering scenario, the contribution from the resonances is greater than $\eta^{{}^{\prime}}$ and $\omega$ combined. The contributions from the resonances are distinct at low $p_{T}$, after 1.5 GeV their contribution vanishes. It is noteworthy that the thermal photons dominate the low $p_{T}$ region and these resonances’ effect can significantly impact their identification. A clear depiction can be seen in Fig 3.
@@
In a rescattering scenario, the extra decay photons can affect the $T_{eff}$ if not removed from the data. We removed the prompt photons and the photons produced in rescattering from the experimental data and fitted the excess photon spectra slope with the exponential function described above to extract the effective temperature (inverse slope $T_{eff}$) of the medium formed. The best match was observed for $T_{eff}$ = 0.167 GeV/c for central Au-Au collisions at 200 GeV. A similar fitting was performed in the corresponding paper [47]. However, the experimental data (which still has the background decay photons ) was used for the fitting, the previous estimate of $T_{eff}$ was about 0.178 GeV/c for central collisions at 200 GeV which is higher than the current values obtained from the fit. A clear depiction of excess photons and the exponential fits can be found in Fig 4.
System $\sqrt{s_{NN}}$ GeV Centrality % $P_{T0}^{ref}$ GeV/c $\frac{dN_{ch}}{d\eta}|_{\eta=0}$
Au-Au 200 (0-20)2.67 519.5
Au-Au 62.4 (0-20)2.85 356.07
Au-Au 39 (0-20)3.40 289.7
Au-Au 27 (0-20)4.10 246.7
d-Au 200 (0-20)3.07 17.8
3He-Au 200 (0-20)3.07 8.7
p-p 200 min bias 4.5 2.38
@@ III.2 Photon production in smaller systems
In Fig 7, we have shown the $\frac{dN_{ch}}{d\eta}^{1.25}$ scaled $p_{T}$ spectra of decay photons for central Au-Au collisions at 200 GeV, 62.4 GeV, 39 GeV, 27 GeV, and for p-p, 3He-Au, and d-Au collisions at 200 GeV. The decay photons show a scaling behavior similar to the ones observed for direct photons in experiments. The scaling continues up to 1 GeV, after 1 GeV it becomes beam energy dependent. It is expected as the increase in multiplicity increases the number of decays. The hard photons show no such trend. The scaled yield of decay photons decreases with decreasing beam energy. At a given collision energy the scaled yield of decay photons in smaller systems like p-p, 3He-Au, etc is greater than the larger systems like Au-Au.
@@ IV Summary
In the current work, we have simulated the transverse momentum spectra of direct as well as decay photons in Au-Au, d-Au, and 3He-Au collisions using the PYTHIA 8/ Angantyr. The hadronic decays produce most of the photons in an Au-Au collision. The main contribution come from the $\pi^{0}$ meson followed by $\eta$, $\eta^{,}$ and $\omega$. The contributions from the rest of the decays are fairly small. The direct photons produced in our simulations are a combination of primary hard scatterings, multi-parton interactions, and initial state radiations. Most of the low $p_{T}$ photons are produced in initial state radiation while the high $p_{T}$ photons are produced in the primary hard scatterings. The Angantyr algorithm divides the sub-collisions into ND, SD, and DD events hence the calculations are lower than a $N_{coll}$ scaled p-p collision. The model shows a good match with the experimental data at high $p_{T}$.

### v2

@@ I Introduction
The major sources of photons in the A-A or p-p collision systems include hard scattering, secondary hard scattering, soft processes, initial state and final state radiations, and resonance decays [12]. The weak showers can also be a source of photons but their effects are very small. Other exotic processes involving one extra dimension can also produce photons via graviton exchange. However, such processes are beyond the scope of this work [13]. The major distinction between a p-p collision and an A-A collision is the existence of QGP which itself emits photons, such photons are called thermal photons. Many theoretical approaches have been made to study direct photons like the elliptic fireball, thermalized glasma, hydrodynamic simulations of the “fireball” evolution, etc [14, 15, 16, 17]. The recent experiments under the RHIC Beam Energy Scan (BES) measured the direct photon spectra for lower collision energies like $\sqrt{s_{NN}}=$ 62.4 GeV and 39.0 GeV. The measurements found the excess of direct photons which are thermal. Thermal photons can also be found in smaller collision systems like d-Au and 3He-Au [18, 19]. These systems show QGP-like signals such as jet quenching and elliptic flow but fail to produce signals like baryon enhancement. The existence of thermal photons in these systems would imply the existence of QGP. However, effectively distinguishing thermal photons from other sources is a difficult problem due to the contribution of a large number of processes. The major aim of this study is to find the contributions from different processes in photon spectra which will help to distinguish thermal photons from others. We are using PYTHIA8 Angantyr for this study [20, 21]. The Angantyr model combines individual NN sub-collisions from PYTHIA8 based on the FRITIOF model. The model does not consider a hot thermalized medium (QGP). Hence, it can give a clear picture of the non-collective photon spectra in various collisions. A detailed description of the model can be found in the next section.
This letter is organized as follows, in section II we have briefly described the Angantyr model. In section III A we have shown different sources of photons and the effect of parton showers, Mulit-Parton interactions, and rescattering. In section III B we have discussed the prompt photon production in smaller collision systems. Finally, in section IV we have summarized our work.
@@ II Angantyr
PYTHIA8 recently developed the Angantyr algorithm to simulate nucleus-nucleus collisions and proton-nucleus collisions [20, 21]. The FRITIOF program and the wounded nucleon model inspired the model with added effects from partonic hard scattering [22]. The positions and the number of the interacting nucleons and binary nucleon-nucleon collisions are calculated using Glauber-model-based eikonal approximation in impact-parameter space and Gribov’s corrections on color fluctuations are implemented to include diffractive excitation [23, 24], which appears due to the fluctuations in the nucleon substructure [20]. The effects of such fluctuations can be found in the corresponding paper [25]. The hard partonic sub-collisions, normalized by nucleon-nucleon sub-collisions, play a major role at high energies. The model separates the individual NN sub-collisions as an absorptive, wounded target, wounded projectile, double diffractive, and elastic interactions. The absorptive interactions are treated as pp-like Non-Diffractive interactions. The scenario where a single target or projectile nucleon is wounded is treated similarly to Single-Diffractive interaction, the double diffractive and elastic sub-collisions are treated accordingly. All subevents are then combined to simulate a p-A or A-A event. This can give a better and more accurate description of a heavy ion event than a scaled p-p event[26].
System $\sqrt{s_{NN}}$ GeV Centrality % $p_{T0}^{ref}$ GeV/c $\frac{dN_{ch}}{d\eta}|_{\eta=0}$ data (PHENIX) model/data
Au-Au 200 (0-20)2.44 519.5 519.02 $\pm$26.25 1.000
Au-Au 62.4 (0-20)2.68 341.4 341.175 $\pm$ 29.325 1.001
Au-Au 39 (0-20)2.85 278.4 278.0 $\pm$ 24.05 1.001
Au-Au 27 (0-20)2.88 242.05 241.0 $\pm$ 21.5 1.004
d-Au 200 (0-20)3.09 16.85 16.8 $\pm$ 1.17 1.002
3He-Au 200 (0-20)3.05 23.0 22.9$\pm$1.6 1.004
@@
Figure 3: The photon spectra in Au-Au collisions at 200 GeV with rescattering.
@@
Where $\sqrt{s_{ref}}$ is the reference center of mass(c.m) energy. The multiplicity of the produced particles is sensitive to $p^{ref}_{T0}$ [30, 27]. A lower value of $p^{ref}_{T0}$ can lead to high multiplicity. The value of $p^{ref}_{T0}$ is tuned for various collision systems and energies tuned by matching the charged particle multiplicity $\frac{dN_{ch}}{d\eta}$ with the experimental data within uncertainties. The values of $p^{ref}_{T0}$ for different collision systems and the corresponding $\frac{dN_{ch}}{d\eta}$ can be found in Table 1. The final hadronization is done using the LUND string fragmentation model and the probability of creating hadrons from the initial state of partons is computed using LUND area law [31]. The channels for prompt photon production used for the work are $gg\rightarrow q\gamma,q\bar{q}\rightarrow g\gamma$ $,gg\rightarrow g\gamma$ $,q\bar{q}\rightarrow\gamma\gamma$ , and $gg\rightarrow\gamma\gamma$, the matrix elements are from the corresponding papers [32, 33, 34]. These processes are also included in the soft part of the model and their production is controlled using $p^{0}_{T}$.
Besides the hard processes, the Quantum Electrodynamics (QED) processes during Parton showers are a major source of photons. The charged particles produce photons in the initial state showers by Bremmstraulung. QED shower in PYTHIA8 incorporates a fully coherent multipole treatment of photon radiation off systems of charged fermions, vectors, and scalar particles, as well as photon splittings to pairs of charged fermions [35, 36, 37, 38, 39].
@@ III.1 Different sources of photons
The photon spectra can be distinguished in two parts: 1. Direct photons and 2. Decay photons. In Fig.1 we have shown the photon spectra for central Au-Au collisions at 200 GeV, the direct and decay photons are shown separately for better visualization. It can be seen that the yield from the decay is much higher in magnitude than that of the prompt photons. The major contributor to the decay photon yield is the neutral pion ($\pi^{0}$) followed by $\eta$ meson, while the contributions from $\eta^{\prime}$ and $\omega$ are small but still greater than the prompt photons. The other particles such as $\Lambda$, $\phi$, etc. which are not removed from the data in the experiments have a little contribution compared to other sources but they exceed the prompt photons at low $p_{T}$ and their contributions should be considered for a more accurate identification of direct photons.
The prompt photons are a combination of primary hard scattering, Multiparton interactions (MPI), and parton showers. The Angantyr simulations with all processes enabled, with MPI=off, and with both MPI=off and ISR=off are shown in Fig 1. This suggests that apart from primary hard scattering photons are produced in the secondary interactions and Bremmstrahulang-like processes. It can be seen that most of the low $p_{T}$ photons are produced from the initial state radiation (ISR) in the early stage of collision. The simulation without MPI has an invariant yield similar to that of the ”MPI on” case. This suggests that MPI does not significantly affect photon production at high $p_{T}$. However, the MPIs can produce photons in the subsequent hard processes (which is absent in HardQCD simulations), enhancing the low $p_{T}$ spectra. Regardless of the configuration the model matches with the experimental data [28] at high $p_{T}$ ($>5$ GeV) within error bars, this is expected as the high $p_{T}$ photons are due to the contribution from primary hard scatterings. The prompt photon yield in Angantyr is lower than a typical scaled p-p collision yield. This is due to sub-collisions being separated into absorptive, wounded target, wounded projectile, double diffractive, and elastic interactions. A clear depiction can be seen in Fig 1.
@@
The processes after the hadronisation like the hadronic rescatterings also produce photons. The rescattering is significantly high in a heavy-ion event compared to a typical p-p event due to its high multiplicity, hence it should be considered while studying the final state particles. The physics processes are similar to low-energy QCD processes [21]. The hard QCD processes and MPIs are not included as the energy of these collisions will be way below 10 GeV. A detailed rescattering framework used in this work can be found in the corresponding paper[42]. The rescattering starts early in Angantyr due to the absence of QGP, hence its effects in Angantyr are more pronounced than the QGP-based models like UrQMD and SMASH which use rescattering as afterburner[43, 44]. Considering the low energy scales of the processes the prompt photon production is unlikely but the same can not be told about the decay photons. Many exotic resonances are formed during the rescatterings and decay into photons (these photons are often considered direct but we will refer to them as decay photons due to the nature of their origin). The major contributors are particles like the $f_{0}(600)$ meson which is produced abundantly in rescattering due to its low mass. In a rescattering scenario, the contribution from the resonances is greater than $\eta^{\prime}$ and $\omega$ combined. The contributions from the resonances are distinct at low $p_{T}$, after 1.5 GeV their contribution vanishes. It is noteworthy that the thermal photons dominate the low $p_{T}$ region and these resonances’ effect can significantly impact their identification. A clear depiction can be seen in Fig 3.
@@
In a rescattering scenario, the extra decay photons can affect the $T_{eff}$ if not removed from the data. We removed the prompt photons and the photons produced in rescattering from the experimental data and fitted the excess photon spectra slope with the exponential function described above to extract the effective temperature (inverse slope $T_{eff}$) of the medium formed. The best match was observed for $T_{eff}$ = 0.167 GeV/c for central Au-Au collisions at 200 GeV. A similar fitting was performed in the corresponding paper [46]. However, the experimental data (which still has the background decay photons ) was used for the fitting, the previous estimate of $T_{eff}$ was about 0.178 GeV/c for central collisions at 200 GeV which is higher than the current values obtained from the fit. A clear depiction of excess photons and the exponential fits can be found in Fig 4. In the presence of a thermalized medium, the effect of rescattering is expected to be lower than Angantyr which will lead to a higher $T_{eff}$ than the one extracted in this work. The $T_{eff}$ should still be lower than the previous estimates which calculated $T_{eff}$ by fitting the data without considering rescattering.
@@ III.2 Photon production in smaller systems
In Fig 7, we have shown the $\frac{dN_{ch}}{d\eta}^{1.25}$ scaled $p_{T}$ spectra of decay photons for central Au-Au collisions at 200 GeV, 62.4 GeV, 39 GeV, 27 GeV, and for p-p, 3He-Au, and d-Au collisions at 200 GeV. The decay photons show a scaling behavior similar to the ones observed for direct photons in experiments. The scaling continues up to 1 GeV, after 1 GeV it becomes beam energy dependent. It is expected as the increase in multiplicity increases the number of decays. The hard photons show no such trend. The scaled yield of decay photons decreases with decreasing beam energy. At a given collision energy the scaled yield of decay photons in smaller systems like p-p, 3He-Au, etc is greater than the larger systems like Au-Au. This suggests that the smaller systems produce more background photons for a given size.
@@ IV Summary
In the current work, we have simulated the transverse momentum spectra of direct as well as decay photons in Au-Au, d-Au, and 3He-Au collisions using the PYTHIA 8/ Angantyr. The hadronic decays produce most of the photons in an Au-Au collision. The main contribution come from the $\pi^{0}$ meson followed by $\eta$, $\eta^{,}$ and $\omega$. The contributions from the rest of the decays are fairly small. The direct photons produced in our simulations are a combination of primary hard scatterings, multi-parton interactions, and initial state radiations. Most of the low $p_{T}$ photons are produced in initial state radiation while the high $p_{T}$ photons are produced in the primary hard scatterings. The Angantyr algorithm divides the sub-collisions into absorptive, wounded target, wounded projectile, double diffractive, and elastic interactions hence the calculations are lower than a $N_{coll}$ scaled p-p collision. The model shows a good match with the experimental data at high $p_{T}$.